OpenStudy (anonymous):

A cylindrical flower vase is 3 inches across the top and 9 inches high. How much water can it hold? a. 63.6 cu. in. b. 254.3 cu. in. c. 169.6 cu. in. d. 127.2 cu. in.

OpenStudy (jdoe0001):

 \(\bf \textit{volume of a cylinder}=\pi r^2 h\quad h=height\quad r=radius=\frac{diameter}{2}\)

OpenStudy (anonymous):

Im still not understanding it..